Use the inner product <u, v>= 2u1v1 + u2v2 in R2 and the Gram-Schmidt orthonormalization process to transform  {(−2, 1), (−2, 7)}

into an orthonormal basis. (Use the vectors in the order in which they are given.)
Answer #1

Given that lot IR? Solhr <u,us = 24,4+ u 2 and {(-2,1),(-2,7)} 04 = (-2,1), de (-2,7) and {B, Bas be orthogonal basis for an出 m <p: A) A op P. Now, (21) (from) (9) (-2, )) 3 and i =(34) B. A. : : IP, <4.A.yjx (A) = 244 + Now, 十 3 3 3 3 32 256 9 9 32